OpenStudy (anonymous):

GCF of 6a^3b, 4a^2b^2, and 22ab^3

OpenStudy (anonymous):

let's see if can we find gcf of 6,4 and 22 it would be 2 now let's put it aside and find gcf of \[a ^{3}b, 4a ^{2}b ^{2},22ab^3 \] by just seeing it we can take ab common so we get gcf equal to 2ab

OpenStudy (anonymous):

Okay smarty pants...lol GCF 3x^4, 3x, -30

OpenStudy (anonymous):

c'mon you should called me something nice ;) ;) no gcf because we have two different kinds of terms here.. constant and variables and for gcf we need terms of only one kind.. ;)!

OpenStudy (anonymous):

So then what would I write in the box provided?

OpenStudy (anonymous):

no gcf.. may be none

OpenStudy (anonymous):

The base of a triangle is 3 cm greater than the height. The area is 14 cm^2. Find the height and the length of the base.

OpenStudy (anonymous):

okie... here we go.. assume height to be x from given info then base is x+3 now we have formula for area as \[A(\Delta)=1/2 \times base \times height\] by putting given values and values of base and height we obtained we get \[14=1/2 \times x \times (x+3)\] multiply both sides by 2 then we get \[28=x(x+3)\rightarrow x ^{2}+3x-28=0\] can you solve this quadratic equation?

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

I'm truly lost :(

OpenStudy (anonymous):

did you understand how we got quadratic equation?

OpenStudy (anonymous):

No :'(

OpenStudy (anonymous):

c'mon what's there not to understand?

OpenStudy (anonymous):

Hmmm...I'm completely lost. Idk if it's because I been looking at these problems ALL DAY :(

OpenStudy (anonymous):

c'mon take some rest and come back later.. may be that's the way..